OpenStudy (anonymous):

Solve the following quadratic trig equations algebraically (without graphing technology) for the domain 0 ≤ x ≤ 360°. Use exact solutions whenever possible. a.6cos2(θ) + cos(θ) - 1 = 0 b.2 - 2sin2(θ) = cos(θ)

OpenStudy (anonymous):

\[6\cos^2\theta + \cos\theta-1=0\\ (3\cos\theta-1)(2\cos\theta+1)=0\\ \cos\theta=-\tfrac{1}{2},\tfrac{1}{3}\] Can you try it from here?